y= -3 cos 2 (x+π/5)


1
Expert's answer
2021-02-28T17:03:02-0500

Given equation is -

"y=-3cos2(x+\dfrac{\pi}{5})"


Putting y=0 in the above equation we get,

"\Rightarrow 0=-3cos(2x+\dfrac{2\pi}{5})"


"\Rightarrow 2x+\dfrac{2\pi}{5}=cos^{-1}0"


"\Rightarrow 2x=1-\dfrac{2\pi}{5}"


"\Rightarrow x=\dfrac{5-5\pi}{10}"


Also putting x=0 in eqs.(1) we get-

"y=-3cos\dfrac{2\pi}{5}"


"=-3(0.3090)=-0.92"


The Graph of the given equation is-